In 1896 , the first Green Jacket Golf Championship was held. The winner's prize money was $200. In 2018, the winner's check was $2,460,000.


a. What was the annual percentage increase in the winner's check over this period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. If the winner's prize increases at the same rate, what will it be in 2050? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 1,234,567.89 .)
Answer #1

FV = PV(1+r)^n

2460000 = 200 (1+r)^ 122

(1+r)^ 122 = 2460000 / 200

= 12300

1+r = 12300^(1/122)

= 1.0802

r = 1.0802 - 1

= 0.0802 i.e 8.02%

Prize Money in 2050:

FV = PV (1+r)^n

= 2460000 (1+0.0802)^32

= 2460000 (1.0802)^32

= 2460000 * 11.8068

= $ 29,044,816.79